# Variational Transformer Ansatz for the Density Operator of Steady States in Dissipative Quantum Many-Body Systems

### v1

@@ Abstract
The transformer architecture, known for capturing long-range dependencies and intricate patterns, has extended beyond natural language processing. Recently, it has attracted significant attention in quantum information and condensed matter physics. In this work, we propose the transformer density operator ansatz for determining the steady states of dissipative quantum many-body systems. By vectorizing the density operator as a many-body state in a doubled Hilbert space, the transformer encodes the amplitude and phase of the state’s coefficients, with its parameters serving as variational variables. Our design preserves translation invariance while leveraging attention mechanisms to capture diverse long-range correlations. We demonstrate the effectiveness of our approach by numerically calculating the steady states of dissipative Ising and Heisenberg spin chain models, showing that our method achieves excellent accuracy in predicting steady states.
@@
The evolution of an open quantum system is governed by the master equation, and several methods have been developed to solve it in recent years. These include analytic approaches based on the Keldysh formalism Sieberer et al. (2016), tensor network techniques such as the density matrix renormalization group and matrix product operator methods Schollwöck (2005); Orus (2018); Cirac et al. (2021); Cui et al. (2015); Kshetrimayum et al. (2017); Werner et al. (2016), the cluster mean-field approach Jin et al. (2016), phase space methods Luo et al. (2022), and corner-space renormalization Finazzi et al. (2015), among others.
Variational methods are fundamental in the study of quantum many-body systems, offering deep insights into the properties of highly complex physical systems. Neural networks have the capacity to efficiently extract hidden patterns from large datasets Bishop et al. (1995); Fausett et al. (1994). In recent years, neural network-based variational ansatz states have garnered significant attention for solving quantum problems, see e.g. Jia et al. (2019a); Carleo et al. (2019); Jia et al. (2019b, 2020); Gao and Duan (2017); Deng et al. (2017); Zhang et al. (2022). The most well-studied examples are Restricted Boltzmann Machine (RBM) states Carleo and Troyer (2017). Beyond RBMs, other architectures, such as deep Boltzmann machines, convolutional neural networks (CNN), and feedforward neural networks have also been employed to construct neural network ansatz states. Many of these neural network ansatz methods have been extended to open quantum systems Torlai and Melko (2018); Yoshioka and Hamazaki (2019); Hartmann and Carleo (2019); Nagy and Savona (2019); Vicentini et al. (2019); Nigro (2021); Mellak et al. (2024), where density operators are encoded into neural networks.
@@
Figure 1: Illustration of a dissipative spin chain with periodic boundary condition and its transformer representation of density operator. The dissipative rate $\gamma$ describes the strength of the coupling to the environment, which leads to decoherence and information loss in the system.
@@
The transformer architecture has recently gained significant attention due to its success in natural language processing tasks Vaswani et al. (2017). It has also been successfully applied to many-body problems in closed quantum systems Zhang and Di Ventra (2023); Viteritti et al. (2023); Rende et al. (2025). However, its application as an ansatz for solving open quantum systems remains relatively unexplored.
@@
In Ref. Luo et al. (2022), the quantum state is mapped to a probability distribution in phase space, and its evolution is reformulated as a probabilistic equation, enabling the transformer to simulate open quantum system dynamics. In this work, we introduce the transformer density operator ansatz, based on the vectorization of the density operator—an approach that has recently gained attention in studies of open-system quantum phases, weak and strong symmetries, tenfold classification, and related topics (see, e.g., Kawabata et al. (2023); Bao et al. (2023); Sohal and Prem (2025); Yoshioka and Hamazaki (2019); Ma and Turzillo (2024)). We employ this ansatz variationally to solve for the steady state of dissipative quantum systems. As we will demonstrate using the dissipative spin chain model, this approach can efficiently capture the steady state with high precision.
@@
Transformer density operator ansatz. — Consider an $N$-particle system with the Hilbert space $\mathcal{H}$ spanned by the basis states $|\bm{\alpha}\rangle$, where $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{N})$ labels the states of the $N$ degrees of freedom composing the system. For example, in a qubit system, $\alpha_{i}$ take values in $\{0,1\}$. For a density operator $\rho\in B(\mathcal{H})$ (where $B(\mathcal{H})$ denotes the space of all linear operators), which is a positive semidefinite, trace-one and Hermitian operator, we can express its matrix elements as $\rho(\bm{\alpha},\bm{\beta})=\langle\bm{\alpha}|\rho|\bm{\beta}\rangle$ in the basis of $|\bm{\alpha}\rangle$ and $|\bm{\beta}\rangle$. By vectorizing, $\rho$ can be transformed into a vector $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=\sum_{\bm{\alpha},\bm{\beta}}\rho(\bm{\alpha},\bm{\beta})|\bm{\alpha}\rangle|\bm{\beta}\rangle$ in the doubled Hilbert space $\mathcal{H}\otimes\mathcal{H}$ (see supplementary material for further details). In order to construct a variational transformer representation of the density operator, we express the vectorized density operator as
@@
Our transformer density operator ansatz is mainly parameterized by incorporating convolutional layers for local feature extraction and the transformer block with a self-attention mechanism to capture long-range correlations within the density matrix structure. Specifically, our ansatz is entirely parametrized by real-valued parameters, and the final complex output is obtained by combining two real-valued outputs that represent its real and imaginary components. For illustration in Fig. 2, we set the batch number to 1 in the schematic, with additional details provided in Sec. II of the supplementary material.
Our goal is to compute $\rho_{\bm{\theta}}(\bm{\alpha},\bm{\beta})$ for each configuration $(\bm{\alpha},\bm{\beta})$. This requires sampling from the configuration space, as detailed in Sec. III.1 of the supplementary material. Below, we provide a step-by-step discussion on obtaining the steady state.
@@
Subsequently, these local feature embeddings are processed by a transformer encoder block Vaswani et al. (2017) with the self-attention module to capture long-range correlations that can emerge globally in a strongly interacting Ising chain. We introduce three learnable matrices $Q$, $K$, and $V$, each of which transforms an input feature vector into a corresponding query, key, or value representation. Concretely, for any feature vector $\bm{x}_{i}$, we define
@@
where $Q$, $K$, and $V$ share the same shape but are learned independently to capture different aspects of the input features. The attention mechanism allows each site to attend to all other sites by computing the learned attention weights using a scaled dot product, followed by a softmax operation
@@
The context vectors $\{\bm{a}_{1},\dots,\bm{a}_{N}\}$ encode global correlations across the entire system. The output context vectors ${\bm{a}_{1},\dots,\bm{a}_{N}}$ are computed in parallel and added with feature vectors $\{\bm{x}_{1},\dots,\bm{x}_{N}\}$. To further enhance the model’s capacity to capture diverse interactions, the self-attention mechanism can be extended to multi-head attention. In this setting, the feature channels are split into $m$ heads, with independent sets of query, key, and value matrices $Q^{\mu}$, $K^{\mu}$, and $V^{\mu}$ (for $\mu=1,\dots,m$) applied to each head; the outputs from all heads are then concatenated to form the final representation.
The attention mechanism enables the network to capture arbitrary pairwise relationships, which is particularly beneficial in open quantum systems where dissipation and quantum coherence can induce correlations of long-range neighbors. What’s more, with the multi-head attention, the ansatz may exhibit multiple distinct correlation stereotypes since different heads can specialize in capturing different scales of correlation.
Figure 2: Schematic representation of the transformer density operator ansatz for the steady-state density operator of an open quantum spin chain with periodic boundary condition. The input spin configurations $(\bm{\alpha},\bm{\beta})$ are split into left and right components and then stacked and reshaped to form the input to two convolutional layers with circular padding, which embed the local features. A self-attention block then captures long-range dependencies by allowing each spin site to attend to all others through learned attention weights. Global average pooling ensures translation invariance, followed by a fully connected layer that maps the spatially averaged vectors into the real and imaginary parts of a complex output. The last step symmetrizes this output to enforce the Hermiticity of the density operator. For the experiments on both Ising and Heisenberg chains presented in this paper, we employ the exact same architecture.
@@
Figure 3: We employ a variational transformer density operator as the steady-state ansatz for a 16-site dissipative transverse-field Ising chain with periodic boundary conditions, uniform dissipative rate, and an interaction strength of $V=2\gamma$. The model is trained using a combination of Stochastic Gradient Descent and the Stochastic Reconfiguration method to optimize the variational parameters. The red points in the figure represent the expectation values $\langle\sigma_{x}\rangle$, $\langle\sigma_{y}\rangle$, and $\langle\sigma_{z}\rangle$, computed from the optimized transformer ansatz, demonstrating its capability to accurately capture the steady-state properties of the system. The exact curve is calculated using NetKet.
Variational algorithm for searching steady state based on transformer density operator ansatz. — Consider a quantum system $\mathcal{H}_{S}$ with $\operatorname{dim}\mathcal{H}_{S}=d$. When coupled to a Markovian environment $\mathcal{H}_{E}$, the evolution equation of the system takes the form of the Gorini–Kossakowski–Sudarshan–Lindblad (GKSL) equation Lindblad (1976); Gorini et al. (1976), also known as the quantum Liouville equation or master equation:
@@
where the Lindbladian $\mathcal{L}$ is a superoperator, $H$ is the Hamiltonian, and $L_{i}$’s are the jump operators associated with the dissipative processes induced by the environment. The $\gamma_{i}$’s are the dissipation rates. There are at most $d^{2}-1$ jump operators over $\mathcal{H}_{S}$. The GKSL equation is the most general equation satisfying the following constraints: (i) local in time, (ii) ensures the positivity $\rho(t)\geq 0$ for all $t$, (iii) is trace-preserving, i.e., $\operatorname{Tr}\hat{\rho}(t)=1$ for all $t$, and (iv) forms a quantum dynamical semigroup.
@@
Since $\operatorname{Loss}(\bm{\theta})\geq 0$. The loss function becomes zero if and only if the steady state is reached. The parameters $\bm{\theta}$ of the transformer that achieve this will give the desired steady state. This ansatz can be modeled by neural network and optimized using variational Monte Carlo methods as introduced in Sec. III.5 in supplementary material. Through variational Monte Carlo optimization, the parameters $\bm{\theta}$ are adjusted so that the resulting density operator accurately represents the steady state of the open quantum system under study. The full optimization procedure is described in detail in Sec. III of the supplementary material.
@@
Figure 5: The observables $\langle\sigma_{x}\rangle$, $\langle\sigma_{y}\rangle$, and $\langle\sigma_{z}\rangle$ are evaluated as functions of the transverse magnetic field $B_{y}/\gamma$ for the case $N=5$, with parameters $J_{x}/\gamma=1.4$, $J_{y}/\gamma=2.0$, $J_{z}/\gamma=1.0$, $B_{x}/\gamma=-1.0$, and $B_{z}/\gamma=0.1$. The exact expectation values (black line) were obtained using exact diagonalization via NetKet, serving as a baseline for comparison. Our transformer density operator ansatz (red dots) achieves excellent agreement across the entire range of $B_{y}/\gamma$, validating its effectiveness in approximating the quantum state and observable dynamics.
@@
In our numerical simulations, we study a system of 16 lattice sites with periodic boundary conditions. The dissipation rate $\gamma_{i}$ in Eq. (8) is taken to be uniform across all sites, and we set the coupling constant to $V=2\gamma$. To obtain the steady-state density matrix, we employ our transformer density operator ansatz, optimizing it using the SGD or Adam optimizer in combination with the stochastic reconfiguration algorithm Chen and Heyl (2024).
@@
To evaluate the accuracy of our ansatz, we compute the expectation values of local observables, specifically the steady-state magnetization components along the $x$, $y$, and $z$ directions, given by $\langle{\sigma}_{k}\rangle_{\mathrm{ss}}=\frac{1}{N}\sum_{i=1}^{N}\langle\sigma^{k}_{i}\rangle,k\in\{x,y,z\}$. These expectation values are estimated via Monte Carlo sampling, following the procedure outlined in Sec. III.2 of the supplementary material. Specifically, we obtain matrix elements of the density operator for a set of sampled configurations, from which we approximate the expectation values using a local estimator approach. A detailed formulation of the method, including the probability distribution used in the sampling process and the construction of the local estimator, can be found in Secs. III.1 and III.2 of the supplementary material.
@@
In Fig. 3, we present $\langle\sigma_{x}\rangle$, $\langle\sigma_{y}\rangle$, and $\langle\sigma_{z}\rangle$ as a function of the normalized transverse field strength $g/\gamma$. The solid black lines correspond to the exact steady-state values, while the red points denote the results obtained using our transformer density operator ansatz. The exact result is obtained using the iterative BiCGStab solver for large systems as described in III.6 of the supplementary material. As shown in Fig. 3, the trained results closely match the exact values. This demonstrates that our ansatz effectively captures the essential physics of the steady state in the dissipative quantum system.
@@
To provide a more comprehensive demonstration of the performance of our ansatz, we also show the optimization curve for a transverse field of magnitude $g=1.6$, as presented in Fig. 4, which illustrates the convergence of the loss function $\operatorname{Loss}(\bm{\theta})=\langle\hat{\mathcal{L}}^{\dagger}\hat{\mathcal{L}}\rangle_{\rho_{\bm{\theta}}}$ as a function of the number of iterations. Detailed optimization strategy is introduced in Sec. III.5.3 of the supplementary material. The transformer density operator ansatz of the blue curve demonstrates strict convergence. This indicates that the transformer density operator ansatz is well equipped to capture correlations inherent in the quantum Ising chain, owing to its self-attention mechanism. Such correlations are essential for accurately modeling the steady-state properties of dissipative quantum systems. This makes the transformer density operator a promising ansatz for addressing systems with complex correlation structures.

### v2

@@ Abstract
The transformer architecture, known for capturing long-range dependencies and intricate patterns, has extended beyond natural language processing. Recently, it has attracted significant attention in quantum information and condensed matter physics. In this work, we propose the transformer density operator ansatz for determining the steady states of dissipative quantum many-body systems. By vectorizing the density operator as a many-body state in a doubled Hilbert space, the transformer encodes the amplitude and phase of the state’s coefficients, with its parameters serving as variational variables. Our design preserves translation invariance while leveraging attention mechanisms to capture diverse long-range correlations. We demonstrate the effectiveness of our approach by numerically calculating the steady states of dissipative Ising and Heisenberg spin chain models, showing that our method achieves excellent accuracy in predicting mixed steady states.
@@
The evolution of an open quantum system is governed by the master equation, and several methods have been developed to solve it in recent years. These include analytic approaches based on the Keldysh formalism Sieberer et al. (2016), tensor network techniques such as the density matrix renormalization group and matrix product operator methods Schollwöck (2005); Orus (2018); Cirac et al. (2021); Cui et al. (2015); Kshetrimayum et al. (2017); Werner et al. (2016), the quantum-trajectory method Mølmer et al. (1993); Vicentini (2019), the cluster mean-field approach Jin et al. (2016), phase space methods Luo et al. (2022), and corner-space renormalization Finazzi et al. (2015), among others.
@@
Figure 1: Illustration of a dissipative spin chain with periodic boundary condition (top) and its transformer representation of the density operator (bottom). The dissipative rate $\gamma$ describes the strength of the coupling to the environment, which leads to decoherence and information loss in the system.
Variational methods are fundamental in the study of quantum many-body systems, offering deep insights into the properties of highly complex physical systems. Neural networks have the capacity to efficiently extract hidden patterns from large datasets Bishop et al. (1995); Fausett et al. (1994). In recent years, neural network-based variational ansatz states have garnered significant attention for solving quantum problems, see e.g. Jia et al. (2019a); Carleo et al. (2019); Jia et al. (2019b, 2020); Gao and Duan (2017); Deng et al. (2017); Zhang et al. (2025). The most well-studied examples are Restricted Boltzmann Machine (RBM) states Carleo and Troyer (2017). Beyond RBMs, other architectures, such as deep Boltzmann machines, convolutional neural networks (CNN), and feedforward neural networks have also been employed to construct neural network ansatz states. Many of these neural network ansatz methods have been extended to open quantum systems Torlai and Melko (2018); Yoshioka and Hamazaki (2019); Hartmann and Carleo (2019); Nagy and Savona (2019); Vicentini et al. (2019); Nigro (2021); Mellak et al. (2024); Kothe and Kirton (2024), where density operators are encoded into neural networks. However, most existing neural quantum state approaches for open systems rely on architectures with inherent locality constraints, limiting their effectiveness in capturing the complex, long-range correlated mixed states that commonly arise as steady states of dissipative systems.
Figure 2: Schematic representation of the transformer density operator ansatz for the steady-state density operator of an open quantum spin chain with periodic boundary conditions. The input spin configurations $(\bm{\alpha},\bm{\beta})$ are split into left and right components and then stacked and reshaped to form the input to two convolutional layers with circular padding, which embed the local features. A self-attention block then captures long-range dependencies by allowing each spin site to attend to all others through learned attention weights. Global average pooling ensures translation invariance, followed by a fully connected layer that maps the spatially averaged vectors into the real and imaginary parts of a complex output. The last step symmetrizes this output to enforce the Hermiticity of the density operator. For the experiments on both Ising and Heisenberg chains presented in this paper, we employ the exact same architecture.
@@
The transformer architecture has recently gained significant attention due to its success in natural language processing tasks Vaswani et al. (2017). It has also been successfully applied to many-body problems in closed quantum systems Zhang and Di Ventra (2023); Viteritti et al. (2023); Rende et al. (2025). However, its application as a steady state ansatz for solving open quantum systems remains relatively underexplored. The key advantage of transformers lies in their self-attention mechanism, which can efficiently capture arbitrary long-range correlations without the locality constraints inherent in tensor network methods. This makes them particularly well-suited for open quantum systems, where dissipation can induce complex, non-local correlation patterns in the steady state.
@@
In Ref. Luo et al. (2022), the quantum state is mapped to a probability distribution in phase space, and its evolution is reformulated as a probabilistic equation, enabling the transformer to simulate open quantum system dynamics. In this work, we introduce the transformer density operator ansatz, based on the vectorization of the density operator—an approach that has recently gained attention in studies of open-system quantum phases, weak and strong symmetries, tenfold classification, and related topics (see, e.g., Kawabata et al. (2023); Bao et al. (2023); Sohal and Prem (2025); Yoshioka and Hamazaki (2019); Ma and Turzillo (2024); Kothe and Kirton (2024)). We employ this ansatz variationally to solve for the steady state of dissipative quantum systems. As we will demonstrate using the dissipative spin chain model, this approach can efficiently capture the steady state with high precision.
@@
Transformer density operator ansatz. — Consider an $N$-particle system with the Hilbert space $\mathcal{H}$ spanned by the basis states $|\bm{\alpha}\rangle$, where $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{N})$ labels the states of the $N$ degrees of freedom composing the system. For example, in a qubit system, $\alpha_{i}$ take values in $\{0,1\}$. For a density operator $\rho\in B(\mathcal{H})$ (where $B(\mathcal{H})$ denotes the space of all linear operators), which is a positive semidefinite, trace-one and Hermitian operator, we can express its matrix elements as $\rho(\bm{\alpha},\bm{\beta})=\langle\bm{\alpha}|\rho|\bm{\beta}\rangle$ in the basis of $|\bm{\alpha}\rangle$ and $|\bm{\beta}\rangle$. By vectorization, $\rho$ can be transformed into a vector $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=\sum_{\bm{\alpha},\bm{\beta}}\rho(\bm{\alpha},\bm{\beta})|\bm{\alpha}\rangle|\bm{\beta}\rangle$ in the doubled Hilbert space $\mathcal{H}\otimes\mathcal{H}$ (see supplementary material for further details). In order to construct a variational transformer representation of the density operator, we express the vectorized density operator as
@@
Our transformer density operator ansatz is mainly parameterized by incorporating convolutional layers for local feature extraction and the transformer block with a self-attention mechanism to capture long-range correlations within the density matrix structure. Specifically, our ansatz is entirely parametrized by real-valued parameters, and the final complex output is obtained by combining two real-valued outputs that represent its real and imaginary components. For illustration in Fig. 2, we set the batch number to 1 in the schematic, with additional details provided in Sec. B of the supplementary material.
Our goal is to compute $\rho_{\bm{\theta}}(\bm{\alpha},\bm{\beta})$ for each configuration $(\bm{\alpha},\bm{\beta})$. This requires sampling from the configuration space, as detailed in Sec. C.1 of the supplementary material. Below, we provide a step-by-step discussion on obtaining the steady state.
@@
Subsequently, these local feature embeddings are processed by a transformer encoder block Vaswani et al. (2017) with the self-attention module to capture long-range correlations that can emerge globally in a strongly interacting Ising chain. We introduce three learnable matrices $Q$, $K$, and $V$, each of which transforms an input feature vector into a corresponding query, key, and value representation. Concretely, for any feature vector $\bm{x}_{i}$, we define
@@
where $Q$, $K$, and $V$ share the same shape but are learned independently to capture different aspects of the input features. The attention mechanism $\omega(\cdot,\cdot)$ allows each site to attend to all other sites by computing the learned attention weights using a scaled dot product, followed by a softmax operation
@@
The context vectors $\{\bm{a}_{1},\dots,\bm{a}_{N}\}$ encode global correlations across the entire system. The output context vectors ${\bm{a}_{1},\dots,\bm{a}_{N}}$ are computed in parallel and added with feature vectors $\{\bm{x}_{1},\dots,\bm{x}_{N}\}$. To further enhance the model’s capacity to capture diverse interactions, the self-attention mechanism is extended to multi-head attention. In this setting, the feature channels are split into $m$ heads, with independent sets of query, key, and value matrices $Q^{\mu}$, $K^{\mu}$, and $V^{\mu}$ (for $\mu=1,\dots,m$) applied to each head; the outputs from all heads are then concatenated to form the final representation.
The attention mechanism enables the network to capture arbitrary pairwise relationships, which is particularly beneficial in open quantum systems where dissipation and quantum coherence can induce correlations of long-range neighbors. Furthermore, with the multi-head attention, the ansatz may exhibit multiple distinct correlation stereotypes since different heads can specialize in capturing different scales of correlation.
@@
Figure 3: We employ a variational transformer density operator as the steady-state ansatz for a 16-site dissipative transverse-field Ising chain with periodic boundary conditions, uniform dissipative rate, and an interaction strength of $V=2\gamma$. The red points in the figure represent the expectation values $\langle\sigma_{x}\rangle$, $\langle\sigma_{y}\rangle$, and $\langle\sigma_{z}\rangle$, computed from the optimized transformer ansatz. These points precisely match the exact solid black line, demonstrating its capability to accurately capture the steady-state properties of the system.
Variational algorithm for searching steady state based on transformer density operator ansatz. — Consider a quantum system $\mathcal{H}_{S}$ with $\operatorname{dim}\mathcal{H}_{S}=d_{S}$. When coupled with a Markovian environment $\mathcal{H}_{E}$, the evolution equation of the system takes the form of the Gorini–Kossakowski–Sudarshan–Lindblad (GKSL) equation Lindblad (1976); Gorini et al. (1976), also known as the quantum Liouville equation or master equation:
@@
where the Lindbladian $\mathcal{L}$ is a superoperator, $H$ is the Hamiltonian, and $L_{i}$’s are the jump operators associated with the dissipative processes induced by the environment. The $\gamma_{i}$’s are the dissipation rates. There are at most $d_{S}^{2}-1$ jump operators over $\mathcal{H}_{S}$. The GKSL equation is the most general equation satisfying the following constraints: (i) local in time, (ii) ensures the positivity $\rho(t)\geq 0$ for all $t$, (iii) is trace-preserving, i.e., $\operatorname{Tr}\hat{\rho}(t)=1$ for all $t$, and (iv) forms a quantum dynamical semigroup.
@@
For the loss function, we have $\operatorname{Loss}(\bm{\theta})\geq 0$, and it reaches zero if and only if the steady state is reached. The parameters $\bm{\theta}$ of the transformer that achieve this will give the desired steady state. This ansatz can be modeled by a neural network and optimized using variational Monte Carlo methods as introduced in Sec. C.5 in the supplementary material. Through variational Monte Carlo optimization, the parameters $\bm{\theta}$ are adjusted so that the resulting density operator accurately represents the steady state of the open quantum system under study. The full optimization procedure is described in detail in Sec. C of the supplementary material.
@@
Figure 5: Purity of the steady state for a 6-site transverse-field Ising chain with periodic boundary conditions, uniform local dissipative rate $\gamma$ and nearest-neighbour coupling $V=2\gamma$. The graph shows that the purity of the steady state decreases with the increase of the transverse magnetic fields and the purity from our transformer ansatz aligns well with the exact curve. This shows that our transformer ansatz has the capability of learning mixed steady states.
@@
In our numerical simulations, we study a system of 16 lattice sites with periodic boundary conditions. The dissipation rate $\gamma_{i}$ in Eq. (8) is taken to be uniform across all sites, and we set the coupling constant to $V=2\gamma$. To obtain the steady-state density matrix, we employ our transformer density operator ansatz, optimizing its parameters using the Stochastic Gradient Descent or Adam Kingma and Ba (2014) optimizer in combination with the stochastic reconfiguration algorithm Chen and Heyl (2024).
Steady states can be either pure or mixed. To verify that our method accurately captures mixed steady states, we computed the purity $\mathrm{Tr}(\rho_{\bm{\theta}}^{2})$ of the optimized transformer-based density operator ansatz, and compared it with the results obtained via exact diagonalization. The purity, which ranges from 0 to 1, quantifies the degree of mixedness: a smaller purity indicates a more mixed state.
When the transverse field is zero ($g=0$), it is straightforward to verify that the pure state $\chi_{\rm SS}=|1\cdots 1\rangle\langle 1\cdots 1|$ is a steady state with purity one. This follows since $[H,\chi_{\rm SS}]=0$ and $\sigma^{-}|1\rangle=0$. As $g$ increases, the steady state gradually becomes mixed due to the system undergoing local energy dissipation. In the limit $g\gg V$, the Hamiltonian reduces to $H=\frac{g}{2}\sum_{i}\sigma^{x}_{i}$, in which case the qubits decouple and evolve independently. Consequently, the steady state of the full system can be obtained as a tensor product of single-qubit steady states (due to the symmetry).
The steady state condition for a single qubit reads
$-i\frac{g}{2}[\sigma^{x},\rho_{\rm SS}^{(1)}]+\gamma\left(\sigma^{-}\rho_{\rm SS}^{(1)}\sigma^{+}-\frac{1}{2}\{\sigma^{+}\sigma^{-},\rho_{\rm SS}^{(1)}\}\right)=0.$ (12)
For simplicity, we set $\gamma=1$. Solving this equation yields
$\rho_{\rm SS}^{(1)}=\begin{pmatrix}\frac{g^{2}}{1+2g^{2}}&\frac{-ig}{1+2g^{2}}\\ \frac{ig}{1+2g^{2}}&\frac{1+g^{2}}{1+2g^{2}}\end{pmatrix}.$ (13)
Therefore, the steady state of the full Ising chain in this limit reads
$\rho_{\rm SS}=\bigotimes_{i=1}^{N}\rho_{\rm SS}^{(1)}.$ (14)
As $g\to\infty$, the steady state approaches the maximally mixed state: $\rho_{\rm SS}\to\mathds{I}/2^{N}$. In this case, the purity of the state is given by
$\mathrm{Tr}(\rho_{\rm SS}^{2})=\frac{1}{2^{N}}.$ (15)
For a long chain with length $N\gg 1$, the purity decays exponentially and tends to zero, indicating a highly mixed steady state.
We examined a 6-site transverse-field Ising model to demonstrate this capability. As shown in Fig. 5, the purity decreases monotonically as the transverse field strength $g$ increases. The solid black curve is calculated via exact diagnolization using qutip Lambert et al. (2024); Johansson et al. (2012) and the red points are the purity calculated from our trained transformer density operator ansatz. At $g/\gamma=4$, the purity drops to approximately 0.01, indicating a highly mixed steady state. This substantial reduction in purity confirms that the steady state is indeed mixed rather than pure, with the degree of mixing increasing with stronger transverse fields. Importantly, our transformer density operator ansatz accurately reproduces the exact purity values across the entire parameter range, demonstrating excellent agreement with the benchmark calculations.
Figure 6: Expectation values of the averaged megnetization components $\langle\sigma_{x}\rangle$, $\langle\sigma_{y}\rangle$, and $\langle\sigma_{z}\rangle$ are evaluated as functions of the transverse magnetic field $B_{y}/\gamma$ for a 5-site Heisenberg spin chain, with parameters $J_{x}/\gamma=1.4$, $J_{y}/\gamma=2.0$, $J_{z}/\gamma=1.0$, $B_{x}/\gamma=-1.0$, $B_{z}/\gamma=0.1$ and periodic boundary conditions. Our transformer density operator ansatz (red dots) achieves excellent agreement across the entire range of $B_{y}/\gamma$, validating its effectiveness in approximating the quantum state and observable dynamics.
@@
To further evaluate the accuracy and expressiveness of our ansatz, we compute the expectation values of local observables, specifically the steady-state magnetization components along the $x$, $y$, and $z$ directions, given by $\langle{\sigma}_{k}\rangle_{\mathrm{ss}}=\frac{1}{N}\sum_{i=1}^{N}\langle\sigma^{k}_{i}\rangle,k\in\{x,y,z\}$, for different magnetic field strengths $g$. These expectation values are estimated via Monte Carlo sampling, following the procedure outlined in Sec. C.2 of the supplementary material. Specifically, we obtain matrix elements of the density operator for a set of sampled configurations, from which we approximate the expectation values using a local estimator approach. A detailed formulation of the method, including the probability distribution used in the sampling process and the construction of the local estimator, can be found in Secs. C.1 and C.2 of the supplementary material.
@@
In Fig. 3, we present $\langle\sigma_{x}\rangle$, $\langle\sigma_{y}\rangle$, and $\langle\sigma_{z}\rangle$ as a function of the normalized transverse field strength $g/\gamma$. The solid black curves represent numerically exact steady-state measurement values obtained by combining the quantum-trajectory algorithm with the MPO method Vicentini et al. (2019); Vicentini (2019). The red points show the predictions of our variational transformer density operator ansatz. Across the entire field range the measurements from our trained ansatz match closely with the black exact curve. The time complexity of the quantum-trajectory method is exponential to the system size, while the computational load of our variantional transformer ansatz grows slowly with additional spins. Therefore our ansatz effectively captures the essential physics of the steady state in the dissipative quantum system.
@@
To provide a more comprehensive demonstration of the performance of our ansatz, we also show the optimization curve for a transverse field of magnitude $g=1.6$, as presented in Fig. 4, which illustrates the convergence of the loss function $\operatorname{Loss}(\bm{\theta})=\langle\hat{\mathcal{L}}^{\dagger}\hat{\mathcal{L}}\rangle_{\rho_{\bm{\theta}}}$ as a function of the number of iterations. Detailed optimization strategy is introduced in Sec. C.5.3 of the supplementary material. The transformer density operator ansatz of the blue curve demonstrates strict convergence. This indicates that the transformer density operator ansatz is well equipped to capture correlations inherent in the quantum Ising chain, owing to its self-attention mechanism. Such correlations are essential for accurately modeling the steady-state properties of dissipative quantum systems. This makes the transformer density operator a promising ansatz for addressing systems with complex correlation structures.